OpenStudy (anonymous):

Hey hey hey peeps :) so I really need FAST! Please answer this question if you are good at algebra, Explain how to simplify the ratio x^2-x-20/x^2+8x+16

hero (hero):

Hint: 1. Factor the numerator and denominator: \[\frac{(x + 4)(x - 5)}{(x + 4)(x + 4)}\] 2. Then cancel factors of one

OpenStudy (anonymous):

You would cancel out the x+4 in the numerator and one of the x+4 in the denominator right?

OpenStudy (anonymous):

Could you explain the WHOLE process please?

OpenStudy (amistre64):

that was the whole process factor, then "cancel" like terms top to bottom

OpenStudy (amistre64):

only thing that might be added is to microwave some pizza rolls ....

OpenStudy (anonymous):

ha ha -_- jk jk:) thanks but seriously I mean explain everything like your talking through the process with a teacher or something

OpenStudy (amistre64):

.... teachers like it when your succinct :) factor the polynomials, and if there are any like factors, they "cancel" out. the term cancel is really not a good phrase to use since you are really using the property that a/a = 1 and 1 times anything is itself. so it just looks like they vanish into thin air

OpenStudy (amistre64):

how wold you go aboutit? that might be a better way to approach this :)

zepdrix (zepdrix):

Monet do you understand how factoring works? :)

OpenStudy (anonymous):

Yes and no..... its kinda confusing.....

hero (hero):

What's confusing about factoring a polynomial and cancelling factors of one?

OpenStudy (anonymous):

Everything :P and I just have a hard time with working out the problem. Also I'm just retarded when it comes to math hee hee :)

OpenStudy (anonymous):

No but for reals I need someone to explain in detail please!

hero (hero):

Anything divided by itself equals one

hero (hero):

Thus (x+4)/(x+4) = 1

hero (hero):

You know how to factor, right?

OpenStudy (anonymous):

But the question is asking you to EXPLAIN how to simplify the ratio x^2-x-20/x^2+8x+16 not write it out

hero (hero):

I gave you the explanation steps as well bro. Look up ↑

OpenStudy (anonymous):

No you said, Anything divided by itself equals one, Thus (x+4)/(x+4) = 1, You know how to factor, right? You did Not explain like you would if you were doing homework bro